OpenStudy (anonymous):

Please help quick!!!! @jim_thompson5910 @ganeshie8 @inkyvoyd @hartnn @wolfe8 @Jamierox4ev3r @abb0t @happinessbreaksbones

OpenStudy (anonymous):

Write the equation for the graphed function. Coordinate Grid A. f(x) = |x – 3| B. f(x) = |x + 3| C. f(x) = |x| – 3 D. f(x) = |x| + 3

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

Anybody? @wolfe8 @happinessbreaksbones

OpenStudy (happinessbreaksbones):

my apologies but I'm horrible at math :( i'm sure wolf8 or someone will get to you shortly.

HanAkoSolo (jamierox4ev3r):

what you could do is graph all these on your calculator...and see which one matches the given graph.

HanAkoSolo (jamierox4ev3r):

lol that's what i would do if you had a graphing calculator :P

OpenStudy (wolfe8):

Sorry I'm not in a good mood tonight but I'll try. Well when you see graphs that are symmetrical along an x value like that, it most probably means you have an absolute value for y. Now one hint: remember the general form of a straight line, y=mx+c

OpenStudy (anonymous):

i dont have a graphing calculator @Jamierox4ev3r and @wolfe8 i dont get it.

OpenStudy (anonymous):

not good right now im sleepy.

OpenStudy (anonymous):

Not being rude but I need answer please. Sorry If im being rude just in a hurry to sleep.

OpenStudy (wolfe8):

Hmm I think you said this last time also. I'm really not into giving answers right away. If you want an answer but don't want explanation, you can graph each one of those online*since you said you don't have a graphing calculator). Just paste the equation in Google's search box.

HanAkoSolo (jamierox4ev3r):

right just graph them all on google or something :P you should have a graphing calc. tho, it would really help you!

OpenStudy (anonymous):

wait when did I say that last time? @wolfe8

OpenStudy (anonymous):

I swear im not lying.

OpenStudy (wolfe8):

Weren't you the one who said he was sleepy and needed answer to go to sleep and was sorry? Also, really, the easiest way to an answer is to graph it online. That's what most people would do too. Plus, it would help you to get familiar with common functions and their graphs.

OpenStudy (anonymous):

ok ok

OpenStudy (wolfe8):

To help you a little, try B and D.